The perimeter of a building is 320 metres the length of the building is 100 meters what is the width?
Answer by checkley77(12844)

P=2L+2W
L=100
320=2*100+2W
320=200+2W
2W=320-200
2W=120
W=60 M. FOR THE WIDTH.
PROOF:
320=2*100+2*60
320=200=120
320=320